OpenStudy (anonymous):

explain how to evaluate log^6 64

Parth (parthkohli):

\( \color{Black}{\Rightarrow \log_6 64}\)?

OpenStudy (anonymous):

yes

Parth (parthkohli):

That wouldn't be a whole number, so I am thinking at the moment.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {\log64 \over \log6} }\) And now you use a calculator.

mathslover (mathslover):

You can simplify log 64 if you want as 2 log 8

Parth (parthkohli):

Yep. Let me explain how you get \(2 \log8\). \( \color{Black}{\Rightarrow \log(8)^2 }\) You bring 2 to the left according to a property. \( \color{Black}{\Rightarrow 2 \log 8}\)

Parth (parthkohli):

Two steps heh

mathslover (mathslover):

\[\large{\log(k^2)=2\log{(k)}}\]

OpenStudy (anonymous):

okay can yu guys help me on my second problem now

Parth (parthkohli):

\(\log_ab^c = c \log_ab\) Better :)

mathslover (mathslover):

*best*